At the time t = 0 s, an object of mass m = 10.5kg
Answer #1

6)

15N

b)

F=ma\\ \implies 15=10.5a\\ \implies 1.43m/s^2=a

c)Along x direction there is no force hence no acceleration, so position is

x=12\times10=120m

Along y direction acceleration is a, so position is

y=\frac{1}{2}\times 1.43\times 10^2=71.5m

So (x,y)=(120,71.5)m

7)

Braking force mg

Let braking force be B

So mg-B=ma\\ \implies mg-ma=B\\ \implies 1.15\times(9.8-8.9)=B\\ \implies 1.035N=B